OpenStudy (anonymous):

Simplify the fraction 10 over the square root of 2 A. five-halves times the square root of 2 B. 5 times the square root of 2 C. 10 times the square root of 2 D. 5

OpenStudy (anonymous):

\[10/\sqrt{2} = 10\sqrt{2}/2=5\sqrt{2}\]

OpenStudy (anonymous):

so answer is B